OpenStudy (anonymous):

Which is a quadratic function? f(x) = 2x + x + 3 f(x) = 0x2 – 4x + 7 f(x) = 5x2 – 4x + 5 f(x) = 3x3 + 2x + 2

OpenStudy (anonymous):

Quadratic function is the form of \[f(x) = ax^2+bx+c~~~\text{where}~~~a>0\]

OpenStudy (anonymous):

@uri @Compassionate

OpenStudy (anonymous):

You know, all you have to do is compare what I said with your options.

OpenStudy (anonymous):

@iambatman May You Help Me If I Need It PLEASE

OpenStudy (anonymous):

Did you figure it out?

OpenStudy (anonymous):

Yes It Was C @iambatman

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

\[f(x) = ax^2+bx+c \implies f(x) = \color{red}5x^2-\color{red}4x+\color{red}7\]